OpenStudy (anonymous):

Find f. f''(x) = −2 + 12x − 12x^2, f(0) = 7, f'(0) = 18

OpenStudy (anonymous):

f ''(x) = −2 + 30x − 12x^2 then using integral f'(x)= −2x + 15x^2 − 4x^3+C1 f'(0)= 0+C1=18 then c1=18 f'(x)= −2x + 15x^2 − 4x^3+18 f(x)=−x^2 + 5x^3 − x^4+18x +C2 f(0)= 0+C2=4 then C2=18 f(x)=−x^2 + 5x^3 − x^4+18x +4 or f(x)=-x^4+5x^3-x^2+18x+4

OpenStudy (anonymous):

thanks! i see what i did wrong now

OpenStudy (anonymous):

your welcome